18:18 overcast, taking darks and flats for last night observations.

22:15 conditions slightly improving, moving to GOODS-N

22:30 seeing probably ok (guider ~1.2") but we are concerned by transparency

22:43 now on A851 (greg), avg guider seeing is ~1.1-1.3"

22:56 starting MOS observations of A851 in J-band

00:00 done with A851 (1h total), observations are not useful, clouds getting thicker.

00:15 guiding is impossible, we are closing

05:00 done for tonight

-- Main.Clement - 06 Mar 2013
Topic revision: r4 - 06 Mar 2013, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback